explain molar mass and molecular mass with there formulas.
1
Expert's answer
2015-07-13T09:29:45-0400
The molecular mass of a substance is the weight in atomic mass units of all the atoms in a given formula. The molar mass is the mass of 1 mole of a substance.
For example:
The molecular mass (Mr) of H2O = 2 · (Ar(H)+Ar(O) = 2 · 1+16 = 18,
where Ar – is the atomic mass;
and the molar mass (M) of H2O = Mr · 1 mole = 18 · 1 = 18 g/moles;
Mr of NaOH = Ar(Na) + Ar(O) + Ar(H) = 23+16+1 = 40;
M of NaOH = Mr(NaOH) · 1mole = 40 · 1 = 40 g/moles;
Mr of H2SO4 = 2 · Ar(H) + Ar(S) + 4 · Ar(O) = 2 · 1+32+4 · 16 = 98
M of H2SO4 = Mr(H2SO4) · 1mole = 98 · 1 = 98 g/moles